# Reproducing NevIR: Negation in Neural Information Retrieval

### v1

@@ 4.2. Datasets and evaluation
We evaluate the different families of model on several IR and negation datasets, namely:
@@
ExcluIR. Unlike the NevIR dataset, which primarily focuses on understanding negation semantics within documents, ExcluIR emphasizes the exclusionary nature of queries. The ExcluIR dataset (Zhang et al., 2024) serves as both an evaluation benchmark and a training set designed to help retrieval models comprehend exclusionary queries. The training set contains 70,293 exclusionary queries generated by ChatGPT-3.5, each paired with a positive and a negative document. The evaluation benchmark consists of 3,452 manually annotated exclusionary queries. For our fine-tuning experiments, we use the benchmark dataset, which is divided into a train set (2,070 pairs, 60%), a validation set (346 pairs, 10%), and a test set (1,040 pairs, 30%).
@@ 4.3. Listwise LLM Re-rankers
Unlike other model categories that rely on triplets for fine-tuning, LLMs require conversational datasets. To facilitate this, we created conversational datasets for both NevIR and ExcluIR and uploaded them to Hugging Face [^8] https://huggingface.co/datasets/thijmennijdam/NevIR-conversational [^9] https://huggingface.co/datasets/thijmennijdam/ExcluIR-conversational. The dataset construction follows the instructional permutation generation method from RankGPT (Sun et al., 2023), enabling models to directly output a ranked list given a set of candidate passages. This approach involves presenting a group of passages, each assigned a unique identifier, to the LLM, which is then prompted to generate their permutation in descending order of relevance to the query.
Sliding Window Strategy. To evaluate the Listwise LLM re-rankers on MS MARCO, we addressed the token limitations of LLMs by adopting a sliding window strategy (Sun et al., 2023). This strategy introduces two hyperparameters: window size and step size. Due to computational constraints, the number of documents that can be re-ranked is limited. To mitigate this, we restricted re-ranking to the top-30 documents retrieved by BM25 (Robertson et al., 2009). As we fine-tune to re-rank only two documents at a time, we observed a decline in ranking performance when handling more than two documents. To prevent this issue from impacting evaluation, we set the sliding window size to 2 and the step size to 1.
@@ 4.4. Experimental Details
Hyperparameters. For fine-tuning the different IR models, we used their respective default hyperparameters for training, as recommended in their original papers (Khattab and Zaharia, 2020; Nogueira et al., 2020; Reimers and Gurevych, 2019). Specifically, the late-interaction model ColBERT uses a default learning rate of 3e-6, the bi-encoder model from Sentence-Transformers uses 2e-5, and the cross-encoder model MonoT5 employs a learning rate of 3e-4. For the Mistral-7B-Instruct-v0.3 model we use a learning rate of 2e-4.
@@ 5. Experimental Results
In this section, we present and describe the results on our reproducibility experiments as well as those of the extension experiments. The results of the reproducibility experiments are presented next to those of the original paper to verify reproducibility.
@@ 5.1. Negation Understanding on SotA Neural Architectures
All models originally evaluated exhibit poor performance, with scores significantly below human levels. This prompts the first research question:
@@ 1. RQ1
Table 1. Model performance comparison on NevIR in terms of pairwise accuracy. We include both newly tested and reproduced models, showing the original and our reproduced scores where available. An ”X” indicates cases where no original score was reported.
Type Data Params Model Name Score Score (ours)
@@
MS MARCO 61M MonoT5 small (Nogueira et al., 2020) 27.7% 27.7%
@@
MS MARCO 223M MonoT5 base (Nogueira et al., 2020) 34.9% 34.9%
MS MARCO 2.85B MonoT5 3B (Nogueira et al., 2020) 50.6% 50.6%
@@
Regarding the bi-encoders, the GritLM-7B model achieves significant improvements (39%), outperforming other models in this category by a substantial margin. In contrast, models such as RepLlama, PromptRetriever variants, and gte-Qwen-2 variants show comparatively minor improvements over previously tested bi-encoders, despite utilizing larger LLM backbones such as Llama-2, Qwen (1.5B and 7B), and Mistral-7B. Notably, GritLM also employs a Mistral-7B backbone and functions as both an embedding and generative model, potentially enhancing its negation understanding compared to traditional embedding models. This strong performance of GritLM suggests that bi-encoders may eventually develop more effective representations for negation understanding. However, further analysis is needed to identify the specific factors contributing to this superior performance.
In the cross-encoder category, the integration of Llama-7B in RankLlama (31.6%) does not improve performance over previous SotA models. The bge models we benchmarked also exhibit only limited improvement over prior evaluations. However, the Jina-reranker-v2-base-multilingual model achieves a pairwise accuracy of 65.2% and significantly outperforms all other cross-encoders tested, despite its smaller model size.
@@
To better understand model limitations, we analyzed errors made by GPT-4o and o3-mini. Our qualitative assessment suggests that their mistakes often occur on more challenging samples. For example, one case required distinguishing between invitations to formal and informal exhibitions, where the key factor was whether the event was officially organized. This task necessitates both an understanding of negation and real-world knowledge about formal events, highlighting the inherent difficulty of achieving a perfect score on the NevIR dataset.
5.2. Results on the Negation Learning Experiments.
Reproducing fine-tuning on NevIR. The results of the reproduced fine-tuning experiment on NevIR are presented in Figure 3. Consistent with the original paper, all models show improvements in pairwise accuracy after fine-tuning, with the bi-encoder and late interaction models reaching around 50%, and the cross-encoder achieving approximately 75%. However, none of the models attain perfect accuracy, indicating potential for further improvement. This led us to consider how listwise LLM re-rankers perform when fine-tuned on on the NevIR distribution, trying to address our second research question:
  1. RQ2
Few-shot LLM-based on NevIR Table 2 shows the results of the few-shot experiments. It can be seen that for the smaller model the performance drops when going from zero-shot to few-shot learning. The model does not understand the provided few-shot negation pattern and is unable of generalising it well to the other examples, probably because of its pre-training and size. GPT-4o is big enough to understand the ”negation context” of these few-shot examples and apply them, Mistral does not understand how the examples relate to the test sample and fails to handle them correctly.
@@ 1. RQ2
Model Name Shots NevIR Score
@@
This partly answers RQ2, fine-tuning on NevIR can significantly enhance performance for a listwise LLM re-ranker, suggesting an improved ability to handle negation. However, this conclusion remains uncertain as there may be other contributing factors such as the model learning a shortcut, e.g. the data distribution to solve the task. This led us to consider the third research question:
@@ 1. RQ3
    How does fine-tuning on other negation datasets improve generalisability on NevIR?
@@ 5.3. Examining Patterns of generalised Negation Comprehension
To test how well the negation comprehension learned from the NevIR dataset generalises to different negation scenarios, we evaluate various IR models trained on NevIR on a different negation dataset, ExcluIR, and vice versa. The results are shown in Figure 4.
For the bi-encoder, late interaction and listwise LLM re-ranker models, the ExcluIR performance does not improve when fine-tuning on the NevIR dataset and only shows an increase when the ExcluIR data is included in the training set. Similarly, the NevIR performance also does not improve when fine-tuned on ExcluIR. This suggests that, for these models, the learned negation comprehension does not transfer effectively to other negation scenarios.
In contrast, for the MonoT5 model, we observe that fine-tuning on the NevIR dataset leads to an improvement in the ExcluIR score as well. A similar trend is observed when fine-tuned on ExcluIR and evaluated on NevIR. This indicates that MonoT5 is capable of generalising its negation comprehension to a different dataset, possibly due it’s large model size and expressive representation of cross-encoders.
Additionally, we test all models on a merged dataset where the training sets of NevIR and ExcluIR are combined. Across all models, this results in the highest or near highest accuracy on both datasets, indicating that training on both datasets simultaneously leads to the best overall performance. This supports the idea that NevIR and ExcluIR may capture distinct aspects of negation that require explicit exposure during training to be effectively learned, answering RQ3.
@@
5.4. Analysis of Training Behavior
@@ 1. RQ4
Figure 3 presents the learning curves for NevIR and MS MARCO during fine-tuning on NevIR. Consistent with findings in the original paper, we observe a decline in MRR, particularly in bi-encoder and cross-encoder models, namely multi-qa-mpnet and MonoT5. Their steeper learning curves show that these models may be learning faster but are also more prone to overfitting. Additionally, we evaluate the Mistral-7B-Instruct-v0.3 listwise re-ranker model on MS MARCO using a sliding window approach. Unlike the other models, Mistral-7B-Instruct-v0.3 maintains a stable MRR throughout fine-tuning, indicating that learning to re-rank on NevIR does not inherently degrade its ability to rank documents on MS MARCO.
To further investigate, Figure 5 illustrates the impact of overfitting by comparing rankings produced by the original multi-qa-mpnet model and its fine-tuned counterpart on NevIR for a specific query. After fine-tuning, the model fails at exact term matching, instead favoring documents featuring other well-known individuals. This suggests that learning a new task overrides fundamental ranking heuristics, possibly because the model shifts toward contrastive learning, emphasizing negation handling over traditional lexical matching.
We find that overfitting effects can be partially mitigated by adopting a more effective early stopping criterion. Rather than selecting the best model based solely on NevIR validation performance, we apply a trade-off criterion that optimizes for both NevIR and MS MARCO performance and select the optimal model based on the average of MRR@10 and pairwise accuracy. Table 3 shows that this trade-off selection preserves NevIR performance while significantly reducing degradation on MS MARCO, mitigating overfitting and addressing part of RQ4.
Figure 5. Rankings on MS MARCO made by multi-qa-mpnet base model and it’s fine-tuned version on NevIR.
@@
Model Selection criterion NevIR score MSMarco MRR@10
@@
MonoT5 base NevIR 71.3% 0.16
MonoT5 base Trade-off 71.6% 0.30
@@ 6. Discussion and Conclusion
Beyond reproducing the results of the original paper, we extend the experiments by evaluating newly developed state-of-the-art ranking models. We observe improvements across all tested categories, with notable advancements in bi-encoders and cross-encoders. Specifically, GritLM, a bi-encoder model, demonstrates substantial improvements, achieving performance comparable to many cross-encoders. Similarly in the cross-encoder category, jina-re-ranker-v2-base-multilingual outperforms MonoT5-3B, despite having a significantly smaller model size.
Also, we tested the latest listwise LLM re-rankers on the NevIR task and found that they achieved the highest performance on NevIR, benefiting significantly from scaling. Furthermore, our results suggest that the reasoning capabilities of LLMs enhance performance, as o3-mini outperforms GPT-4o considerably. However, these models come at a significant computational cost, limiting their practicality to final-stage re-ranking.
Additionally, we fine-tune Mistral-7B-Instruct-v0.3 and observe improved performance on NevIR. However, unlike GPT-4o, Mistral-7B-Instruct-v0.3’s performance degrades in few-shot learning, where GPT-4o has substantial gains. This suggests that only GPT-4o is large enough to effectively grasp context and improve negation understanding.
To investigate the generalisation of negation comprehension learned from NevIR, we fine-tuned the models on NevIR and evaluated them on ExcluIR, and vice versa. The results showed that only the cross-encoder model demonstrated effective knowledge transfer of negation comprehension to this new setting. This is likely due to the cross-encoder’s large model size and its ability to combine the query and documents into the most expressive representation among the different model categories. This suggests that the cross-encoder model is the only model that meaningfully represents negation in a way that generalises across different data distributions and contexts.
Despite the positive trend we observe on the NevIR benchmark, our findings show that many neural IR models still struggle with negation. Even the best models continue to fall short of human-level performance, reinforcing that negation understanding remains a persistent challenge for neural IR systems.
@@
One limitation of our assessment of re-ranking with the Mistral model is that we trained it on triplets from NevIR rather than on a list of documents, as NevIR does not contain listwise training data. As a result, we could not evaluate performance degradation on MS MARCO with a sliding window larger than two, since fine-tuning the model on lists longer than two made it unable to correctly handle the re-ranking task.
@@
The total GPU usage for this research, including exploratory runs and repeated experiments, was approximately 100 hours. All experiments were conducted on NVIDIA H100 PCIe GPUs (TDP of 350W) using the Snellius supercomputer, which has a carbon efficiency of 0.421 kgCO2eq/kWh.
We approximated the environmental impact using the MachineLearning Impact calculator with an A100 PCIe 40/80GB GPU (Lacoste et al., 2019). Total emissions were estimated post hoc to be approximately 13.16 kgCO2eq. These estimations were conducted using the MachineLearning Impact calculator presented in (Lacoste et al., 2019) and account for both GPU power consumption and the energy efficiency of the hardware infrastructure.
For experiments with the ChatGPT models, we used the OpenAI API. Running these models requires significant computational resources and water consumption. While exact figures are not available, we highlight the environmental impact and encourage future research to consider environmental responsibility when conducting scientific studies.

### v3

@@ 4.2. Datasets and evaluation
We evaluate the different families of models on several IR and negation datasets, namely:
@@
ExcluIR. Unlike the NevIR dataset, which primarily focuses on understanding negation semantics within documents, ExcluIR emphasises the exclusionary nature of queries. The ExcluIR dataset (Zhang et al., 2024) serves as both an evaluation benchmark and a training set designed to help retrieval models comprehend exclusionary queries. The training set contains 70,293 exclusionary queries generated by ChatGPT-3.5, each paired with a positive and a negative document. The evaluation benchmark consists of 3,452 manually annotated exclusionary queries. For our fine-tuning experiments, we use the benchmark dataset, which is divided into a train set (2,070 pairs, 60%), a validation set (346 pairs, 10%), and a test set (1,040 pairs, 30%).
@@ 4.3. Listwise LLM Re-rankers
Unlike other model categories that rely on triplets for fine-tuning, LLMs require conversational datasets. To facilitate this, we converted NevIR and ExcluIR to datasets with a conversational structure and uploaded them to Hugging Face [^8] https://huggingface.co/datasets/thijmennijdam/NevIR-conversational [^9] https://huggingface.co/datasets/thijmennijdam/ExcluIR-conversational. The dataset construction follows the instructional permutation generation method from RankGPT (Sun et al., 2023), enabling models to directly output a ranked list given a set of candidate passages. This approach involves presenting a group of passages, each assigned a unique identifier, to the LLM, which is then prompted to generate their permutation in descending order of relevance to the query.
Sliding Window Strategy. To evaluate the Listwise LLM re-rankers on MS MARCO, we addressed the token limitations of LLMs by adopting a sliding window strategy (Sun et al., 2023). This strategy introduces two hyperparameters: window size and step size. Due to computational constraints, the number of documents that can be re-ranked is limited. Therefore, we restricted re-ranking to the top-30 documents retrieved by BM25 (Robertson et al., 2009). A problem we encountered was that fine-tuning to re-rank more than two documents at a time (e.g., setting the window size higher than two) on the MS MARCO dataset led to a decline in ranking performance when re-ranking only the two documents in NevIR. To prevent this issue from impacting evaluation on NevIR, we set the sliding window size to 2 and the step size to 1 in all our fine-tuning experiments.
@@ 4.4. Experimental Details
Hyperparameters. For fine-tuning the different IR models, we used their respective default hyperparameters for training, as recommended in their original papers (Khattab and Zaharia, 2020; Nogueira et al., 2020; Reimers and Gurevych, 2019). Specifically, the ColBERT models use a default learning rate of 3e-6, the bi-encoder model from Sentence-Transformers uses 2e-5, and the cross-encoder model MonoT5-base employs a learning rate of 3e-4. For the Mistral-7B-Instruct-v0.3 model, we use a learning rate of 2e-4.
@@ 5. Experimental Results
In this section, we present and describe the results of our reproducibility experiments as well as those of the extensions.
@@ 5.1. Negation Understanding on SotA Neural Architectures
All models originally evaluated exhibit poor performance, with scores significantly below the human baseline of 100% (Weller et al., 2024b). This prompts the first research question:
@@ 1. RQ1
Table 1. Model performance comparison on NevIR in terms of pairwise accuracy (P. Acc). We include both newly tested and reproduced models, showing the original and our reproduced scores where available. An ”X” indicates cases where no original score was reported. Training dataset indicates which dataset the models are trained on from scratch.
Type Training Dataset Params Model Name P. Acc P. Acc (ours)
@@
MS MARCO 61M MonoT5-small (Nogueira et al., 2020) 27.7% 27.7%
@@
MS MARCO 223M MonoT5-base (Nogueira et al., 2020) 34.9% 34.9%
MS MARCO 2.85B MonoT5-3B (Nogueira et al., 2020) 50.6% 50.6%
@@
Regarding the bi-encoders, the GritLM-7B model achieves significant improvements (39%), outperforming other models in this category by a substantial margin. In contrast, models such as RepLlama, PromptRetriever variants, and gte-Qwen-2 variants show comparatively minor improvements over previously tested bi-encoders, despite utilising larger LLM backbones such as Llama-2, Qwen (1.5B and 7B), and Mistral-7B. Notably, GritLM also employs a Mistral-7B backbone and functions as both an embedding and generative model, potentially enhancing its negation understanding compared to traditional embedding models. This strong performance of GritLM suggests that bi-encoders may eventually develop more effective representations for negation understanding. However, further analysis is needed to identify the specific factors contributing to this superior performance.
In the cross-encoder category, the integration of Llama-7B in RankLlama (31.6%) does not improve performance over previous SotA models. The bge-rerankers we benchmarked also show only limited improvement compared to earlier evaluations. In contrast, the Jina-reranker-v2-base-multilingual model achieves a pairwise accuracy of 65.2% and significantly outperforms all other cross-encoders tested, despite its smaller model size. Since both the bge-rerankers and the jina-reranker are trained on multilingual data, we hypothesise that this performance difference is due to the jina-reranker having been trained on a larger amount or higher quality negation data.
@@
Model Name Shots Pairwise Accuracy
@@
We analysed errors made by the top-performing models; GPT-4o and o3-mini. This qualitative assessment suggests that their mistakes often occur on more challenging samples. For example, both made a mistake when the query asks whether the event was officially organised, and the documents distinguish between invitations to formal and informal exhibitions. This task not only necessitates an understanding of negation but also of real-world knowledge about formal events, highlighting the inherent difficulty of achieving a perfect score on the NevIR dataset.
5.2. Results on the Negation Learning Experiments.
Reproducing fine-tuning on NevIR. The results of the reproduced fine-tuning experiment on NevIR are presented in Figure 3. Consistent with the original paper, all models show improvements in pairwise accuracy after fine-tuning, with the bi-encoder and late interaction models reaching around 50%, and the cross-encoder achieving approximately 75%. However, none of the models attain perfect accuracy, indicating potential for further improvement. This led us to consider how listwise LLM re-rankers perform when fine-tuned on NevIR data, prompting our second research question:
  1. RQ2
Few-shot listwise LLM-reranker on NevIR. Table 2 shows the results of the few-shot experiments. Notably, the performance of Mistral-7B-Instruct-v0.3 drops when going from zero-shot to few-shot learning. This suggests that the model struggles to internalise and generalise the provided few-shot negation patterns. In contrast, GPT-4o demonstrates improved performance with the inclusion of more examples, indicating a positive trend in few-shot learning capabilities. These findings align with previous research indicating that larger LMs tend to perform better in few-shot learning scenarios (Brown et al., 2020).
This partly answers RQ2, fine-tuning on NevIR can significantly enhance performance for a listwise LLM re-ranker, suggesting an improved ability to handle negation. However, this conclusion remains uncertain as there may be other contributing factors, such as the model learning a shortcut, e.g. the data distribution, to solve the task. This led us to consider the third research question:
@@ 5.3. Examining Patterns of Generalised Negation Comprehension
To test how well negation comprehension learned from the NevIR dataset generalises to different negation scenarios, we fine-tune various IR models on NevIR and evaluate on a different negation dataset, ExcluIR. To assess generalisation fully, we also fine-tune models on ExcluIR and evaluate on NevIR. The results are presented in Figure 4.
For the bi-encoder, late interaction and listwise LLM re-ranker models, the ExcluIR performance does not improve when fine-tuned on the NevIR dataset and only shows an increase when the ExcluIR data is included in the training set. Similarly, the NevIR performance also does not improve when fine-tuned on ExcluIR. This shows that, for these models, the seeming negation comprehension does not transfer effectively to other negation scenarios.
In contrast, for the MonoT5-base model, we observe that fine-tuning on the NevIR dataset leads to an improvement in the ExcluIR score as well. A similar trend is observed when fine-tuning on ExcluIR and evaluating on NevIR. This indicates that MonoT5-base is capable of generalising its negation comprehension to a different dataset, due to its large model size and the expressive representation of cross-encoders.
Additionally, we test all models on a merged dataset where the training sets of NevIR and ExcluIR are combined. Across all models, this results in the highest or near-highest accuracy on both datasets, indicating that training on both datasets simultaneously leads to the best overall performance.
We answer RQ3 by finding that NevIR and ExcluIR may capture distinct aspects of negation that require explicit exposure during training to be effectively learned. As a result, models show limited to no generalisability across these datasets.
@@
5.4. Analysis of Training Behaviour
@@ 1. RQ4
Figure 3 presents the learning curves for NevIR and MS MARCO during fine-tuning on NevIR. Consistent with findings in the original paper, we observe a decline in MRR, particularly in bi-encoder and cross-encoder models, namely multi-qa-mpnet-base-dot-v1 and MonoT5-base. Their steeper learning curves show that these models may be learning faster, but are also more prone to overfitting. Additionally, we evaluate the Mistral-7B-Instruct-v0.3 listwise re-ranker model on MS MARCO using a sliding window approach. Unlike the other models, Mistral-7B-Instruct-v0.3 maintains a stable MRR throughout fine-tuning, indicating that learning to re-rank on NevIR does not inherently degrade its ability to rank documents on MS MARCO.
To further investigate the impact of overfitting, we compare the ranking produced by the original multi-qa-mpnet-base-dot-v1 model and the variant fine-tuned on NevIR. We observed that after fine-tuning, the model fails at exact term matching, instead favouring documents featuring other well-known individuals. An example highlighting this is provided in Figure 5. This suggests that learning a new task overrides fundamental ranking heuristics, possibly because the model shifts toward contrastive learning, emphasising negation handling over traditional lexical matching.
Figure 5. A query from MS MARCO along with rankings made by multi-qa-mpnet base model and it’s fine-tuned version on NevIR.
@@
Model Selection criterion NevIR P. Acc MSMarco MRR@10
@@
MonoT5-base NevIR 71.3% 0.16
MonoT5-base Trade-off 71.6% 0.30
We find that the effects of overfitting can be partially mitigated by adopting a more effective early stopping criterion. Rather than selecting the best model based solely on NevIR validation performance, we apply a trade-off criterion that optimises for both NevIR and MS MARCO performance and select the optimal model based on the average of MRR@10 and pairwise accuracy. Table 3 shows that this trade-off selection preserves NevIR performance while significantly reducing degradation on MS MARCO, thus mitigating overfitting.
We answer RQ4 by finding that overfitting to negation data can degrade general ranking, but a trade-off-based selection criterion helps maintain performance on both negation and MS MARCO.
@@ 6. Discussion and Conclusion
Beyond reproducing the results of the original paper, we extend the experiments by evaluating newly developed SotA ranking models. We observe improvements across all tested categories, with notable advancements in bi-encoders and cross-encoders. Specifically, GritLM, a bi-encoder model, demonstrates substantial improvements, achieving performance comparable to many cross-encoders. Similarly, in the cross-encoder category, jina-re-ranker-v2-base-multilingual outperforms MonoT5-3B, despite having a significantly smaller model size.
Also, we tested the latest listwise LLM re-rankers on the NevIR task and found that they achieved the highest performance on NevIR, benefiting significantly from scaling. Furthermore, our results suggest that the reasoning capabilities of LLMs enhance performance, as o3-mini outperforms GPT-4o considerably. However, these models come at a significant computational cost, limiting their practicality to final-stage re-ranking. Additionally, we fine-tune Mistral-7B-Instruct-v0.3 and observe improved performance on NevIR. However, unlike GPT-4o, Mistral-7B-Instruct-v0.3’s performance degrades in few-shot learning, where GPT-4o has substantial gains. This indicates that only GPT-4o is large enough to effectively grasp context and improve negation understanding.
To investigate the generalisation of negation comprehension learned from NevIR, we fine-tuned the models on NevIR and evaluated them on ExcluIR, and vice versa. The results show that only the cross-encoder model successfully transferred its negation understanding to the new setting. We attribute this to the cross-encoder’s larger model size and its ability to combine the query and documents into a more expressive joint representation compared to other model categories. We find that the cross-encoder model captures negation in a way that meaningfully generalises across different data distributions and contexts.
Despite the positive trend we observe on the NevIR benchmark, our overall findings show that many neural IR models still struggle with negation. Even the best models continue to fall short of human-level performance, reinforcing that negation understanding remains a persistent challenge for neural IR systems.
@@
One limitation of our assessment of re-ranking with the Mistral model is that we trained it on triplets from NevIR rather than on a list of documents, as NevIR does not contain listwise training data. As a result, we could not evaluate performance degradation on MS MARCO with a sliding window larger than two, since fine-tuning the model on lists longer than two made it unable to correctly handle the re-ranking task on NevIR.
@@
The total GPU usage for this research, including exploratory and repeated runs, was approximately 100 hours on NVIDIA H100 PCIe GPUs (TDP: 350W), using the Snellius supercomputer (carbon efficiency: 0.421 kgCO2eq/kWh).
We estimated emissions post hoc using the MachineLearning Impact calculator (Lacoste et al., 2019), selecting an A100 PCIe 40/80GB GPU. Total emissions were approximately 13.16 kgCO2eq, accounting for GPU power use and infrastructure efficiency.
Experiments using ChatGPT models were conducted via the OpenAI API. These models require substantial computing and water resources. Although exact figures are unavailable, we highlight their environmental cost and encourage future research to account for such impacts.
Acknowledgements.
This research was partly supported by the Swiss National Science Foundation (SNSF), under the project PACINO (Personality And Conversational INformatiOn Access), grant number 215742.